### v1

# CALRec: Contrastive Alignment of Generative LLMs For Sequential Recommendation

Abstract.
Traditional recommender systems such as matrix factorization methods rely on learning a shared dense embedding space to represent both items and user preferences. Sequence models such as RNN, GRUs, and, recently, Transformers have also excelled in the task of sequential recommendation. This task requires understanding the sequential structure present in users’ historical interactions to predict the next item they may like. Building upon the success of Large Language Models (LLMs) in a variety of tasks, researchers have recently explored using LLMs that are pretrained on vast corpora of text for sequential recommendation. To use LLMs in sequential recommendations, both the history of user interactions and the model’s prediction of the next item are expressed in text form. We propose CALRec, a two-stage LLM finetuning framework that finetunes a pretrained LLM in a two-tower fashion using a mixture of two contrastive losses and a language modeling loss: the LLM is first finetuned on a data mixture from multiple domains followed by another round of target domain finetuning. Our model significantly outperforms many state-of-the-art baselines ($+37\%$ in Recall@1 and $+24\%$ in NDCG@10) and systematic ablation studies reveal that (i) both stages of finetuning are crucial, and, when combined, we achieve improved performance, and (ii) contrastive alignment is effective among the target domains explored in our experiments.
Sequential Recommendation, Large Language Models, Contrastive Learning
copyright: noneccs: Information systems Recommender systemsccs: Computing methodologies Neural networks
1. Introduction and Motivation
Recommender systems aim to understand users and items and use the learned knowledge to recommend relevant items for future user interactions. Traditional collaborative filtering methods (Goldberg et al., 1992; Sarwar et al., 2001; Koren et al., 2009) map item and user IDs into a common latent embedding space learned from individual user-item interactions, but these embeddings are often static and fail to capture the dynamic nature of user interests. Recent exploration in sequential recommender employs RNN (Hidasi et al., 2016), CNN (Yan et al., 2019) and the transformer (Kang and McAuley, 2018; Sun et al., 2019) to directly learn the time-evolution of user behaviors. The transformer architecture  (Vaswani et al., 2017; Devlin et al., 2019) stands out for its efficient capability of modeling sequences through the self-attention mechanism, which learns the correlation or causal relationships between different tokens and their influence on future tokens. In the context of recommender systems (RecSys), each input position (token) corresponds to a single user activity, typically a user’s interaction (e.g., purchase) with an item  (Kang and McAuley, 2018; Sun et al., 2019; Zhang et al., 2019; Zhou et al., 2020; Hou et al., 2022). Some approaches rely on item IDs, each represented by a unique token, and model a user’s interaction history as a sequence of item ID tokens (Kang and McAuley, 2018; Sun et al., 2019); other efforts combine more sophisticated features, such as the text embedding of item descriptions, to capture richer content information for each item  (Zhang et al., 2019; Zhou et al., 2020).
Motivated by the global paradigm shift towards autoregressive Language Large Models (LLMs) (Achiam et al., 2023; Anil et al., 2023; Tay et al., 2023; Team et al., 2023) with successful applications in various tasks and domains (Dosovitskiy et al., 2021; Arnab et al., 2021; Ramesh et al., 2022; Yu et al., 2021), we propose using generative LLMs as an approach for sequential recommender systems. Pretrained LLMs learn knowledge from large text corpora and are expected to perform reasonably well in general tasks, and performance improves on specific tasks after finetuning with target domain data. Several studies have explored utilizing LLMs in recommender systems (Wu et al., 2023), such as framing sequential recommendation as sentence-completion (Li et al., 2023c), question-answering (Zhang et al., 2023) or ranking problems (Kang et al., 2023).
Despite the impressive language understanding capabilities of pretrained LLMs, they require fine-tuning for highly specific tasks like sequential recommendation. An LLM should be ideally fully fine-tuned (Kang et al., 2023; Whitehouse et al., 2023) to learn to comprehend domain-specific flattened text sequences and the underlying user interest evolution. Similar to most sequence modeling approaches like SASRec (Kang and McAuley, 2018), BERT4REC (Sun et al., 2019), Recformer (Li et al., 2023b), and GPT4Rec (Li et al., 2023c), we employ next-item prediction as the sequential recommendation task. Based on sentence-completion (Li et al., 2023c) or question-answering  (Zhang et al., 2023) styled prompting, our template design is further inspired by the idea of few-shot learning (Brown et al., 2020). Specifically, we prefix user sequence and each item in the user sequence with distinctive text indicators that 1) allow the model to better understand data pattern and text format and 2) increase the LLM response coherence of the input sentence.
We propose the Contrastive Aligned Generative LLM Recommendation (CALRec) framework to adapt an LLM for sequential recommendation, inspired by the effectiveness of contrastive learning (CL) in mapping-based RecSys (Hou et al., 2022; Li et al., 2023b) and CL research in NLP and other areas  (Gao et al., 2021; Li et al., 2023a). CALRec features the following:
  • Pure text input and text output with advanced prompt design.
  • A mixed training objective that combines customized next-item generation objective and auxiliary contrastive objectives.
  • A two-stage LLM fine-tuning paradigm consisting of a multi-category joint fine-tuning stage followed by a category-specific fine-tuning stage.
  • A novel quasi-round-robin BM25 retrieval approach for item retrieval.
Similar to previous work (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2023b), we leverage the Amazon Review dataset (Ni et al., 2019) and compare our results to open-source state-of-the-art (SotA) baselines. We observed a non-negligible percentage of consecutive identical activities by the same users in the raw datasets which might trivialize the next-item prediction task. We thus train and evaluate the model on deduplicated sequences and propose to adopt ’Last Item Repeater’, a training-free, text-statistic baseline for model comparison (§4 and §6). We also revisit the existing evaluation metrics of the previous works, and propose to report both optimistic and pessimistic metric scores for text-based RecSys (§4). Through comprehensive experiments, we discover that our CALRec method demonstrates superior performance and outperforms existing SotA by a considerable margin for most evaluation metrics. A systematic ablation study indicates the effectiveness of the two-stage training paradigm and our contrastive alignment among the various data categories examined in our experiments.
Finally, we briefly summarize the contribution of our study:
  • We proposed CALRec, a novel sequential recommendation framework that features advanced prompt design, a two-stage training paradigm, a combined training objective, and a quasi-round-robin BM25 retrieval approach.
  • We conduct comprehensive experiments, ablation study and further analyses and show that 1) our approach outperforms existing baselines in most metrics by a considerable margin; 2) the main components of our approach are effective. We also discuss 3) the characteristics of our output and our limitations.
  • We revisit data preprocessing and evaluation metrics in this field.
2. Related Work
Traditional recommender systems like Collaborative Filtering (Goldberg et al., 1992) learn relevance through user-item co-occurrence. Early collaborative filtering approaches like Matrix Factorization (MF) (Koren et al., 2009) are ID-only where both user and item IDs are embedded in a shared embedding space. Later MF improvements incorporated other attributes (Shi et al., 2014) such as time and location. Nevertheless, these approaches did not capture the sequential nature of user behavior which gave rise to a class of Sequential Recommenders. Similar to MF, initial Sequential Recommenders were ID-only like GRU4Rec (Hidasi et al., 2016), SASRec (Kang and McAuley, 2018), and Bert4Rec (Sun et al., 2019), while follow-up work focused on incorporating additional attributes beyond the IDs, with methods such as FDSA (Zhang et al., 2019), S3-Rec (Zhou et al., 2020), and UniSRec (Hou et al., 2022). Due to the challenges (Shi et al., 2014) that IDs pose in real world Recommender Systems, researchers(Hou et al., 2022; Li et al., 2023b) explored removing them in favor of attribute-only models but faced quality challenges because of the valuable memorization characteristics that IDs bring in. Instead, other researchers explored replacing them with IDs derived from attribute embeddings e.g. (Rajput et al., 2023; Singh et al., 2023; Hou et al., 2023) which alleviate some of the infrastructure challenges posed by random IDs in exchange for increased system complexity.
With the rise of pretrained LLM backbones, which are either encoder-only like BERT (Devlin et al., 2019; Li et al., 2023b), encoder-decoder like T5 (Raffel et al., 2020), or decoder-only like PaLM  (Chowdhery et al., 2023; Anil et al., 2023), GPT (Brown et al., 2020), LLama (Touvron et al., 2023), among others, researchers focused on ways to leverage these in the Sequential Recommender tasks primarily to bring along the world knowledge that these models encode in their parameters. Some of these approaches are zero-shot or few-shot (Wang and Lim, 2023; Liu et al., 2023; Hou et al., 2024) where no LLM parameters are updated, others used the LLM to extract item text embeddings also without parameter updates (Hou et al., 2022).
Researchers explored directly finetuning encoder-only, encoder-decoder or decoder-only LLMs to incorporate the Recommender System knowledge into the LLM (Li et al., 2023b; Raffel et al., 2020; Wu et al., 2023; Kang et al., 2023; Zhang et al., 2023; Bao et al., 2023). P5 (Geng et al., 2022) finetuned a pre-trained T5 (Raffel et al., 2020) model to solve 5 common RecSys tasks; Sequential Recommendation is one of them. Unlike P5 which models the Sequential Recommendation task as an ID generation task and its prompt templates does not distinguish between user and item, our work focuses on generating item attribute text and explicitly finetunes the model in a two-tower fashion to explicitly understand the items and their attributes as entries within a user-item interaction sequence. GPT4Rec (Li et al., 2023c) finetunes a pre-trained GPT-2 model to generate the next item titles. Like GPT4Rec, we use BM25 (Robertson et al., 2009) to search for items in the corpus. However, unlike GPT4Rec which directly does supervised finetuning for next item generation for each target domain as a sentence completion task, our work is the first to explore using contrastive loss as training objective for multi-stage finetuning of pre-trained autoregressive language models on sequential recommendation task. This enables the model to generalize to multiple target domains in the final stage of finetuning. We include the list of baseline models we compare against and their details in §4.
3. Methodology
3.1. Preliminaries and Terminology
We assume a set of users $\mathcal{U}$, where a user $u\in\mathcal{U}$ has an interaction sequence that consists of a sequence of $n$ items $(I_{1},I_{2},...,I_{n})$ in chronological order ($n$ can be different for different users). The sequential recommendation task is defined as follows: given the user’s interaction sequence $I_{1:n-1}=(I_{1},I_{2},...,I_{n-1})$, a sequential recommender aims to predict the target item $I_{n}$, the last item in the user’s history. In our study, we use pure text to represent each item and user interaction sequence (§3.2), and we adapt autoregressive LLMs pretrained on large corpora of text for sequential recommendation. Specifically, we pose the task as a sequence-to-sequence generation task where the input is the text description of $I_{1:n-1}$ and the expected output is that of $I_{n}$. The underlying motivation to pose the problem fully in the text domain is to unlock the rich text information in real-world data and directly leverage LLMs’ strong language understanding and reasoning capabilities.
Figure 1. CALRec: Training Framework.
3.2. Data Format and Template Design
[…]
  [USER PURCHASE HISTORY SEQUENCE]
This is the summary of a user’s purchase history. The first item bought is as follows. Title: The Hillman Group 591520 Small Wire Nail and Brad Assortment, 260-Pack. Category: Tools &\&& Home Improvement. Brand: Hillman. Price: Unknown.
The next item bought is as follows. Title: Pro Flex Super Flexible Elastomeric Sealant Acrylic Clear Paintable 10 Oz. Category: Tools &\&& Home Improvement. Brand: Geocel. Price: \$$\$11.22$.
The next item bought is as follows.
  [TARGET ITEM SEQUENCE]
Title: Trimaco Llc RF36 35-Inch by 166-Feet Rosin Flooring Paper, Red. Category: Tools &\&& Home Improvement. Brand: Trimaco. Price: Unknown.
Different from our method, many previous text-based approaches do not specify an item prefix and do not leverage one-off user history suffix (Li et al., 2023b, c; Zhang et al., 2023).
3.3. Two-Stage Fine-Tuning
Previous work (Hou et al., 2022; Li et al., 2023b) indicates that pretraining a RecSys on a large volume of data (various data categories) provides favorable inductive bias for downstream tasks (a specific and smaller-scale data category for evaluation after further fine-tuning on them). Analogously, based on an LLM already pretrained on large, diverse text corpora, we propose a two-stage fine-tuning framework that consists of Multi-Category Joint Fine-Tuning (Stage I) and Category-Specific Fine-Tuning (Stage II): the latter benefits from the former as a result of transfer learning. Our Stage I fine-tuning allows the model to adapt to the sequential recommendation problem setting and learn data patterns in a category-agnostic fashion. Inspired by Xue et al. (2021), we adopt a sampling rate for each data category proportional to $|\mathcal{U}|^{0.3}$, where $|\mathcal{U}|$ is the size (number of users) of each category, for the purpose of combating the data imbalance issue, preventing a few extra-large categories from dominating the Stage I finetuning set. After deriving the multi-category fine-tuned model, we further adapt the model to each category respectively via Category-Specific Fine-Tuning (Stage II). In both stages, the training objective is the same, introduced in §3.4.
3.4. Training Objectives
Next Item Generation Objective. The primary objective we use for LLM finetuning is the Next Item (Text) Generation (NIG) Objective. NIG aims to generate the text description of the target item given the text description of the previous items in the user’s history. Denote text-tokenized user sequence as $(t_{1},t_{2},...,t_{l})$, where $l$ is the tokenized sequence length. The first $m$ tokens belong to all the items in the user sequence except for the last item, followed by $l-m$ tokens for the target item. We propose a next item generation objective to adapt an LLM for sequential recommendation:
(1) $\vspace{-5pt}\mathcal{L}_{NIG}=-\mathbb{E}\!\sum_{j=m+1}^{l}\text{log}\,P(t_{j}|t_{1:j-1};\theta)$
where $\theta$ is the whole set of trainable parameters of the chosen LLM.
Auxiliary Contrastive Alignment. One shortcoming of the NIG, is that it deals with the input and output on the token level rather than on the item/user level. Therefore, in addition to the token-level NIG objective, we also investigate auxiliary contrastive objectives that operate on the item/user level. Specifically, we adopt a two-tower training framework where one tower only takes in the target item, and the mean-pooled feature in the final hidden layer is denoted as $\mathbf{v}^{T}$. The second tower takes the whole user sequence and similarly we denote the mean-pooled feature corresponding to the user history sequence as $\mathbf{v}^{U}$ and the feature for the target item conditioned on the user history as $\mathbf{v}^{T|U}$. We experiment with two contrastive losses for user- and item-level alignments following the InfoNCE loss  (Oord et al., 2018), a standard and robust choice in contrastive learning research  (Chen et al., 2020; Gao et al., 2021; Li et al., 2022, 2023a):
(2) $\begin{split}\mathcal{L}_{TT}=&-\frac{1}{N_{b}}\!\sum_{i=1}^{N_{b}}\text{log}\frac{\text{exp}(\text{cos}(\mathbf{v}^{T|U}_{i},\mathbf{v}^{T}_{i})/\tau_{c})}{\sum_{j=1}^{N_{b}}\text{exp}(\text{cos}(\mathbf{v}^{T|U}_{j},\mathbf{v}^{T}_{i})/\tau_{c})},\end{split}$
(3) $\begin{split}\mathcal{L}_{UT}=&-\frac{1}{N_{b}}\!\sum_{i=1}^{N_{b}}\text{log}\frac{\text{exp}(\text{cos}(\mathbf{v}^{U}_{i},\mathbf{v}^{T}_{i})/\tau_{c})}{\sum_{j=1}^{N_{b}}\text{exp}(\text{cos}(\mathbf{v}^{U}_{j},\mathbf{v}^{T}_{i})/\tau_{c})},\end{split}$
where in-batch negative samples are adopted, $N_{b}$ is the batch size (number of user sequences in a training batch), $\text{cos}(\cdot,\cdot)$ calculates the cosine similarity between two input vectors, and $\tau_{c}$ is the temperature for contrastive alignments.
Our final training objective is a mixture of $\mathcal{L}_{NIG}$ and contrastive losses as shown in Fig 1:
(4) $\begin{split}\mathcal{L}_{CALRec}=(1-\alpha-\beta)\mathcal{L}_{NIG}+\alpha\mathcal{L}_{TT}+\beta\mathcal{L}_{UT}.\end{split}$
3.5. Quasi-Round-Robin BM25 Retrieval
At inference, we prompt the model with the text description of user’s history of interactions. To obtain more than one candidate next item prediction, we conduct temperature sampling (Hinton et al., 2015) to generate $N_{\text{gen}}$ pieces of text as $N_{\text{gen}}$ predictions for the next item. Each of the $N_{\text{gen}}$ predictions is associated with its sequence score $s$ (i.e., log probability, which is a negative real number typically in the range of $(-10,0)$, as observed on our data).
1
Input : $samples$: Output texts sampled from the LLM (size = $N_{\text{gen}}$ )
2
Input : $logprobs$: LLM prediction scores for $samples$ (size = $N_{\text{gen}}$)
3
Input : $C$: Item corpus (size = $N_{c}$)
Input : $N_{\text{preds}}$: a hyperparam controlling the number of generated text samples to use for BM25 matching.
Input : $\epsilon$: a hyperparam used to scale the LLM log probs score before combining with the BM25 score.
Output : The sorted list of item predictions.
4
// Sort $samples$ by score and remove duplicates
5 sorted__\__samples, sorted__\__logprobs $\leftarrow$ SortDescending((samples, logprobs), by=logprobs)
6
7sorted__\__samples, sorted__\__logprobs $\leftarrow$ RemoveDuplicates((sorted__\__samples, sorted__\__logprobs))
8
// Keep the top $N_{\text{preds}}$ samples
9
[…]
12 match__\__scores $\leftarrow$ EmptyMatrix($N_{c}$, $N_{\text{preds}}$)
13 for $i\leftarrow 1$ to $N_{\text{preds}}$ do
       // Find BM25 match scores of entire corpus for each samples
14       bm25__\__scores $\leftarrow$ BM25(top__\__samples[i], C)
15       bm25__\__scores__\__scaled $\leftarrow$ LinearScale(bm25__\__scores, axis=0)
       // Modulate by sequence prediction score.
[…]
17       match__\__scores[:,i] $\leftarrow$ modulated__\__bm25__\__scores
18
candidate__\__scores $\leftarrow$ MaxPool(match__\__scores, axis=1)
  // size = Nc
return ArgSortDescending(candidate__\__scores)
Algorithm 1 Quasi Round Robin BM25 Selection
We remove duplicate generations, sort them by sequence scores in descend order, and retain at most top $N_{\text{preds}}$ unique predictions. Then, in order to match these generated texts with items from our corpus, we rely on BM25 retrieval (Robertson et al., 2009) to calculate the matching scores between each text prediction and the whole item corpus (with size $N_{c}$): this will result in BM25 matching scores organized into a matrix of shape ($N_{c}$, $N_{\text{preds}}$). In order to rank the candidate items, we devise the following heuristics which aims to select the closest matches according to both the LLM scores and BM25 scores in a quasi round robin fashion along the different $N_{\text{preds}}$ text generations. First, we linearly scale each column of the matrix respectively to the range $[0,1]$ to make BM25 scores derived from different LLM predictions comparable. We then penalize each prediction’s BM25 scores (each column) by multiplying the BM25 scores by a factor of $e^{\epsilon s}$, where $s$ is the sequence score associated with the text prediction, and $\epsilon$ is a small positive constant so that $e^{\epsilon s}$ is smaller than but very close to $1.0$. This is to further modulate the BM25 text-based retrieval scores by the LLM generative scores. After that, we max-pool along the $N_{\text{preds}}$ axis and finally derive modulated BM25 matching scores in a single vector of size $N_{c}$, with information from all $N_{\text{preds}}$ recommendations fused in it (we can then derive the ranking of the ground-truth target item in the item corpus for calculating evaluation metric scores, or return top-K recommendations for users), as presented in Algorithm 1222In our preliminary investigation, we found that 1) max-pooling significantly outperforms min-pooling and mean-pooling and 2) there is obvious performance drop if we remove modulation since only max-pooling combined with modulation can make sure that top-1 LLM outputs’ top-1 recommendation is still the top-1 recommendation in the merged list. 3) The overall results are not sensitive to the value of $\epsilon$ as long as $\epsilon$ is small.. Here, we provide another view into the round-robin nature of our approach: given that the scaled BM25 score of each best-match item from the non-duplicate $N_{\text{preds}}$ texts generated is $1$, when they are further modulated (scaled down) using $e^{\epsilon s}$, the final retrieval algorithm roughly follows a round-robin fashion, where the top-1 BM25 retrieval result of each of the generated text are in turn grouped as the final top $N_{\text{preds}}$ retrieval results, when $\epsilon$ is close to $0.0$. Since there are ties, we name our method ’quasi-’ rather than ’strict-’ round-robin.
4. Experimental Setup
Before Dedup After Dedup
Dataset # of Users % of duplicate Purchases # of Items Total Purchase Items/User Purchases/Item Density Words/Item Word Vocab Size Avg. Word Freq
Scientific 9461 5.5% 5282 66644 7.04 12.62 0.0013 22.82 19178 6.3
Instruments 25577 4.0% 10599 214526 8.39 20.24 0.00079 18.43 22256 8.8
Arts 47197 9.4% 22828 411449 8.72 18.02 0.00038 21.38 40342 12.1
Office 44736 (50%) 6.3% 27482 352151 7.87 12.81 0.00029 21.74 56687 10.5
Games 50940 4.7% 17383 457060 8.97 26.29 0.00051 16.39 17087 16.7
Pet 43135 (20%) 5.9% 37712 380623 8.82 10.09 0.00023 19.87 47339 15.8
Table 1. Statistics of Amazon Reviews Dataset.
LLM Selection. Our work leverages PaLM-2 (Anil et al., 2023; Tay et al., 2023), a state-of-the-art LLM pretrained with a mixture of different objectives and demonstrates superb generalization capabilities on a variety of language tasks. Among PaLM-2 variations, we choose PaLM-2 XXS as the backbone for our tasks for trade-off in model capability and inference latency. Larger PaLM-2 size-S does not show a significant advantage against XXS without fine-tuning (see §5). Besides, the scale of our dataset (significantly smaller than text corpora for pretraining LLMs) may lead to overfitting when finetuning larger models (Kirkpatrick et al., 2016; Mitchell et al., 2022).
[…]
User Sequence Deduplication. We have found that the Amazon Review datasets include a non-negligible percentage of consecutive duplicate events in user sequences, where the duplication is identified as two consecutive activities by the same user that have the same item ID (ASIN number), review text, rating, and the same review timestamp. As indicated in Table 1, around $4\%$ to $9.4\%$ of purchased items are exact duplicates of the previous ones in their user sequences. We conduct deduplication (dedup) on all user sequences following our strict definition of duplication above, while we still keep two consecutive purchases of the same item, if they occurred at different time (i.e., they come with different timestamps) or come with different ratings/review text since they are likely the users’ real actions. We note that some previous work did not conduct user sequence deduplication (Li et al., 2023b), which may lead a recommender system to learn a trivial strategy of always recommending the last item in the input sequence.
Implementation Details. We train our CALRec based on XXS-sized PaLM-2 on a cluster of TPUv4 chips (Jouppi et al., 2023); our code is implemented with JAX (Bradbury et al., 2018) [^4] https://github.com/google/jaxand PAX [^5] https://github.com/google/paxml. We adopt a training batch size of $512$ users, a learning rate of $1$e44-4- 4, a maximum input length of $1,024$ tokens, and a maximum output decoding length of $80$ tokens. As a data augmentation trick, during training, we randomly truncate the last $k$ items ($k\in\{0,1,2,3,4\}$) from the full user record (with $n$ items) and always use the last item after truncation (i.e., the $(n-k)$-th item in the original user record) as the target item. We train CALRec with $\alpha=0.125,\beta=-0.025$ for up to $150,000$ steps in Stage I and adopt $\tau_{c}=0.5$ (see Eqs. 23) as the config combination yields best validation scores. We adopt the checkpoint with $135,000$ training steps in Stage I selected on the validation data category, and then further fine-tune the model for Stage II. There are two exceptions: Scientific [^6] We pick the Stage I model at $45,000$ training steps since ‘Scientific’ is significantly smaller than other categories and its own validation set indicates that overfitting occurs at $135,000$ steps and then continue with Stage II training., and Games [^7] For Games category, we adopt $\alpha=0.125,\beta=0$ for both stages and its Stage I checkpoint at $105,000$ steps is selected using the Stage I validation set. It is because we found on the ‘Games’ own validation set, the new setup outperforms the previous setup . All our model selection is based on the NDCG@10 metric following previous work (specifically we adopt its pessimistic estimation as defined in the last paragraph of this section). For each case, the same training objective is used in Stages I and II. At inference, the temperature for decoding is $0.5$, $N_{gen}=32$, $N_{pred}=10$ and $\epsilon=1/5000$. There are two hyper-parameters inherent to BM25: we adopt the default values from the original BM25 API, $k_{1}=1.5$ and $b=0.75$, throughout all our experiments and did not further tune them since we found that the model performance is not sensitive to them.
Baselines. The following are all the baseline approaches we considered. SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) are pure ID-Based approaches, Recformer (Li et al., 2023b) and the ‘Last Item Repeater’ (LIR) are purely text-based, and FDSA (Zhang et al., 2019) and S3-Rec (Zhou et al., 2020) leverage both ID and text. UniSRec (Hou et al., 2022) has two variants and we report its performance in both text-only mode and ID+text mode. More details about each baseline are provided in what follows.  [^8] In our experiments, LIR is implemented on our own, Recformer follows its official implementation. The rest baselines are all provided by the official code repo of UniSRec which is based on RecBole (Zhao et al., 2021): the attribute embeddings of FDSA and S3-Rec are text features from our strongest baseline UniSRec.
  • SASRec (Kang and McAuley, 2018) is a causal sequential recommendation approach implemented with an encoder-only unidirectional transformer. It takes in a sequence of item IDs and predicts the ID embedding of the next item in the last position.
  • BERT4Rec (Sun et al., 2019) adopts a bidirectional transformer architecture for sequential recommendation. Its input consists of a sequence of item IDs and it is trained with the masked language modeling objective. At inference, a ‘mask’ token is appended to the end of the sequence to predict the next item.
  • FDSA (Zhang et al., 2019) applies two self-attentional blocks to address ID sequence and attribute feature sequence, respectively, concatenates their final representations and fuses them for next item prediction.
  • S3-Rec (Zhou et al., 2020) maximizes the mutual information between ID and attribute embeddings via contrastive objectives and it is implemented with an encoder-only transformer network. It consists of a pretraining stage and a fine-tuning stage, both conducted on domain-specific data.
  • UniSRec (Hou et al., 2022) separately embeds each item’s text via a pretrained and fixed BERT cascaded by a trainable adapter. Given a sequence of item embeddings, a sequence representation is then modeled with another small transformer network. We adopt the officially released UniSRec model pretrained on multiple domains and fine-tune it on specific Amazon categories, respectively. We report two fine-tuning setups of UniSRec: (1) text-only setup, (2) ID+text setup where an item embedding is the sum of its text and ID representations.
  • Recformer (Li et al., 2023b) relies on the encoder-only Longformer architecture (Beltagy et al., 2020), token type and item position embeddings are proposed to facilitate learning item and user representations. It also features a pretraining stage on multiple categories with a combination of standard Masked Language Modeling (MLM) loss and a contrastive objective aligning user history and target item representations. We adopt Recformer’s officially released pretrained checkpoint and fine-tune it on specific categories, respectively.
  • Last Item Repeater, or LIR hereafter, denotes a simple yet effective baseline strategy which always recommends the last item in the input sequence as our training-free, text-statistic baseline. Based on the text of the last item, we rely on BM25 to rank the items in the full item collection.
Evaluation Metrics. Following standard practices from prior work, we report NDCG@10, Recall@K (K$\in\{1,10\}$) and MRR@$\infty$ scores. Different from previous work, we report both ‘optimistic’ (opt) and ‘pessimistic’ (pes) calculations of the metrics: when there is a tie such that the ground-truth target item and other candidate items have the same matching score, the ground-truth item ranks first in (opt) and ranks last in (pes) calculation, respectively. It is because 1) there are distinct items (with different item IDs) with the same text description and 2) in rare cases, we found that BM25 may yield identical scores for very similar text pieces. We still treat items in case 1) as distinct items because in real production environments vendors/content creators may intentionally or unintentionally duplicate items. Of course, this will not cause any issue for ID-based approaches due to distinct item IDs, where (opt) and (pes) scores are the same. In the official implementations of our text-based baselines, Li et al. (2023b) only calculates (opt) metric scores which may result in unfair comparisons with ID-based baselines.  Hou et al. (2022) applies “torch.topk()” to derive exactly top $K$ item ids for calculating Recall/NDCG@K: if the $K$-th and the $K+1$-th items are having the same score, only one of them will be included in the output list. So its metric score is a number between our(opt) and (pes) estimates. [^9] To provide more insights on our (opt)/(pes) metrics, readers could consider an extreme scenario where the model output is empty and thus when using BM25 retrieval the matching score is 00 for all candidate items. In this case, the optimistic Recall@K scores will always be $1.0$, while the pessimistic scores will always be $0.0$. To provide a fair assessment of the model’s ability to uniquely identify candidate items, we have thus decided to report both (opt) and (pes) scores. All the evaluations are performed on user-sequence-deduplicated datasets.
5. Experimental Results
ID-Only Methods ID+Text Methods Text-Only Methods
Dataset Metric   (opt)/(pes) BERT4Rec SASRec FDSA S3-Rec UniSRec UniSRec RecFormer LIR CALRec Improv. (opt)
Scientific NDCG@10 0.0411 0.0555 0.0594 0.0414 0.0644 0.0620/0.0580 0.0639/0.0588 0.0565/0.0486 0.0788/0.0740 22.4%
Recall@1 0.0239 0.0079 0.0357 0.0172 0.0199 0.0242/0.0140 0.0298/0.0168 0.0142/0.0052 0.0511/0.0389 43.1%
Recall@10 0.0636 0.1065 0.0887 0.0727 0.1181 0.1101/0.1100 0.1058/0.1057 0.1005/0.0933 0.1124/0.1117 -4.8%
MRR 0.0389 0.0456 0.0555 0.0380 0.0555 0.0546/0.0492 0.0570/0.0502 0.0461/0.0385 0.0730/0.0667 28.1%
Instruments NDCG@10 0.0680 0.0623 0.0796 0.0606 0.0709 0.0661/0.0655 0.0596/0.0585 0.0344/0.0338 0.0909/0.0905 14.2%
Recall@1 0.0494 0.0158 0.0574 0.0202 0.0237 0.0251/0.0236 0.0266/0.0239 0.0040/0.0031 0.0718/0.0706 25.1%
Recall@10 0.0915 0.1130 0.1092 0.1048 0.1255 0.1141/0.1141 0.0940/0.0940 0.0696/0.0692 0.1158/0.1158 -7.7%
MRR 0.0658 0.0528 0.0765 0.0526 0.0613 0.0576/0.0569 0.0535/0.0521 0.0263/0.0257 0.0864/0.0857 12.9%
Arts NDCG@10 0.0547 0.0619 0.0726 0.0602 0.0729 0.0630/0.0624 0.0662/0.0653 0.0443/0.0431 0.0864/0.0853 18.5%
Recall@1 0.0347 0.0147 0.0460 0.0216 0.0213 0.0220/0.0208 0.0279/0.0254 0.0067/0.0051 0.0636/0.0610 38.3%
Recall@10 0.0799 0.1118 0.1061 0.1052 0.1320 0.1103/0.1102 0.1095/0.1095 0.0859/0.0850 0.1140/0.1139 -13.6%
MRR 0.0513 0.0519 0.0680 0.0520 0.0615 0.0546/0.0540 0.0582/0.0569 0.0344/0.0331 0.0815/0.0801 19.9%
Office NDCG@10 0.0545 0.0602 0.0749 0.0607 0.0713 0.0625/0.0616 0.0687/0.0676 0.0391/0.0379 0.0976/0.0966 30.3%
Recall@1 0.0403 0.0134 0.0547 0.0291 0.0213 0.0250/0.0228 0.0328/0.0302 0.0056/0.0040 0.0761/0.0734 39.1%
Recall@10 0.0709 0.1033 0.0982 0.0941 0.1220 0.1023/0.1022 0.1063/0.1063 0.0735/0.0728 0.1213/0.1213 -0.6%
MRR 0.0520 0.0499 0.0710 0.0535 0.0597 0.0545/0.0533 0.0603/0.0588 0.0309/0.0296 0.0925/0.0911 30.3%
Games NDCG@10 0.0334 0.0451 0.0526 0.0419 0.0501 0.0408/0.0407 0.0377/0.0360 0.0211/0.0197 0.0595/0.0585 13.1%
Recall@1 0.0112 0.0039 0.0210 0.0080 0.0075 0.0111/0.0107 0.0126/0.0091 0.0015/0.0009 0.0295/0.0277 40.5%
Recall@10 0.0646 0.0989 0.0957 0.0884 0.1074 0.0820/0.0820 0.0724/0.0723 0.0442/0.0427 0.0986/0.0984 -8.2%
MRR 0.0313 0.0380 0.0483 0.0367 0.0427 0.0372/0.0370 0.0349/0.0326 0.0171/0.0159 0.0510/0.0498 5.6%
Pet NDCG@10 0.0376 0.0448 0.0537 0.0392 0.0574 0.0539/0.0464 0.0590/0.0471 0.0391/0.0354 0.0736/0.0693 24.8%
Recall@1 0.0276 0.0093 0.0394 0.0162 0.0202 0.0295/0.0129 0.0377/0.0130 0.0094/0.0030 0.0570/0.0465 44.7%
Recall@10 0.0499 0.0773 0.0710 0.0647 0.0970 0.0829/0.0826 0.0826/0.0818 0.0680/0.0673 0.0937/0.0934 -3.4%
MRR 0.0365 0.0382 0.0522 0.0349 0.0503 0.0498/0.0399 0.0549/0.0393 0.0317/0.0272 0.0696/0.0639 26.8%
Average ($6$ Categories) NDCG@10 0.0482 0.0550 0.0655 0.0507 0.0645 0.0581/0.0558 0.0592/0.0556 0.0391/0.0364 0.0811/0.0790 23.8%
Recall@1 0.0312 0.0108 0.0424 0.0187 0.0190 0.0228/0.0175 0.0279/0.0197 0.0069/0.0036 0.0582/0.0530 37.3%
Recall@10 0.0701 0.1018 0.0948 0.0883 0.1170 0.1003/0.1002 0.0951/0.0949 0.0736/0.0717 0.1093/0.1091 -6.6%
MRR 0.0460 0.0461 0.0619 0.0446 0.0552 0.0514/0.0484 0.0531/0.0483 0.0311/0.0283 0.0757/0.0729 22.3%
Table 2. Evaluation results of CALRec in comparison with different ID/Text-based sequential recommendation baselines (cf. §4). For text-only methods, we report both optimisitc (opt) and pessimistic (pes) as “(opt)/(pes)”. For ID-based methods, (opt) and (pes) scores are the same. Bold: the highest (opt) scores. Underline: the strongest (pes) scores. We also report the relative improvement (opt) comparing CALRec against (opt) baseline SotA marked with Blue Background per data category per metric.
Scientific Instruments Games
Model Variant NDCG@10 Recall@1 Recall@10 MRR NDCG@10 Recall@1 Recall@10 MRR NDCG@10 Recall@1 Recall@10 MRR
CALRec 0.0788 / 0.0740 0.0511 / 0.0389 0.1124 / 0.1117 0.0730 / 0.0667 0.0909 / 0.0905 0.0718 / 0.0706 0.1158 / 0.1158 0.0864 / 0.0857 0.0595 / 0.0585 0.0295 / 0.0277 0.0986 / 0.0984 0.0510 / 0.0498
- Multi-Category Joint Fine-Tuning (A) 0.0727 / 0.0675 0.04408 / 0.0314 0.1070 / 0.1061 0.0670 / 0.0602 0.0870 / 0.0862 0.0673 / 0.0658 0.1119 / 0.1115 0.0822 / 0.0812 0.0583 / 0.0570 0.0286 / 0.0265 0.0966 / 0.0959 0.0501 / 0.0486
- Category-Specific Fine-Tuning (B) 0.0717 / 0.0670 0.0424 / 0.0300 0.1077 / 0.1073 0.0653 / 0.0588 0.0867 / 0.0862 0.0715 / 0.0702 0.1048 / 0.1048 0.0840 / 0.0834 0.0539 / 0.0530 0.0257 / 0.0238 0.0904 / 0.0902 0.0463 / 0.0451
- Contrastive Alignment 0.0771 / 0.0724 0.0495 / 0.0376 0.1109 / 0.1105 0.0712 / 0.0649 0.0905 / 0.0899 0.0711 / 0.0696 0.1154 / 0.1153 0.0858 / 0.0850 0.0593 / 0.0584 0.0289 / 0.0272 0.0987 / 0.0985 0.0507 / 0.0495
Pretrained PaLM-2 (W/O Fine-Tuning) 0.0326 / 0.0291 0.0090 / 0.0055 0.0661 / 0.0623 0.0267 / 0.0234 0.0257 / 0.0254 0.0041 / 0.0037 0.0585 / 0.0584 0.0197 / 0.0193 0.0150 / 0.0141 0.0031 / 0.0024 0.0325 / 0.0314 0.0134 / 0.0127
Table 3. Ablation result showing the benefits of multi-stage finetuning and constrastive alignment to improve the performance of CALRec. Bold: the highest optimistic (opt) scores. Underline: the strongest pessimistic (pes) scores.
Metric W/ Contr. W/O Contr. Improv. p-value
NDCG@10 0.0811/0.0790 0.0803/0.0782 1.00%/1.02% 0.0076/0.0064
Recall@1 0.0582/0.0530 0.0574/0.0521 1.39%/1.73% 0.0027/0.0004
Recall@10 0.1093/0.1091 0.1084/0.1082 0.83%/0.83% 0.0178/0.019
MRR 0.0757/0.0729 0.0748/0.0720 1.20%/1.25% 0.0066/0.0037
Table 4. Average scores (opt)/(pes) on all $6$ data categories: w/ contrastive alignment vs. w/o contrastive alignment.
Pretrained PaLM-2 W/O Fine-Tuning
Dataset Metric   (opt)/(pes) XXXS XXS S
Scientific NDCG@10 0.0299/0.0252 0.0326/0.0291 0.0321/0.0306
Recall@1 0.0089/0.0034 0.0090/0.0055 0.0082/0.0058
Recall@10 0.0592/0.0534 0.0661/0.0623 0.0660/0.0647
MRR 0.0246/0.0204 0.0267/0.0234 0.0262/0.0247
Instruments NDCG@10 0.0215/0.0210 0.0257/0.0254 0.0234/0.0231
Recall@1 0.0029/0.0022 0.0041/0.0037 0.0043/0.0039
Recall@10 0.0480/0.0476 0.0585/0.0584 0.0512/0.0511
MRR 0.0162/0.0156 0.0197/0.0193 0.0188/0.0185
Table 5. Performance results of pretrained PaLM-2 without any fine-tuning for Sequential Recommendation. Bold: the highest optimistic (opt) scores. Underline: the strongest pessimistic (pes) scores.
Main Results. Table 2 demonstrates that the sequential recommendation capabilities of the proposed CALRec framework are stronger than our baselines. On the $6$ Amazon Review test categories, our model outperforms all $8$ baselines significantly in terms of NDCG@10, Recall@1 and MRR. In terms of Recall@10, our approach outperforms all Text-Only and ID-Only models across the board by a large margin, but slightly underperforms our strongest baseline the ID+Text variant of UniSRec. For our method and $3$ Text-Only baselines, the results in Table 2 verified that there is a gap between optimistic and pessimistic metric scores, demonstrating the necessity of reporting both. Meanwhile, it is worth noticing that our pessimistic scores clearly surpass the optimistic scores of the $3$ text-only models on all $6$ data categories. In addition to the absolute scores, we also calculate the percentage gains comparing CALRec’s optimistic scores against the baseline SotA per each data category and per evaluation metric (different method for different data category and metric). The results are presented in the last column of Table 2. We found that, CALRec is especially strong in Recall@1, outperforming baseline SotA by $37.3\%$ on average; CALRec outperforms baseline SotA by $23.8\%$ and $22.3\%$ on average in terms of NDCG@10 and MRR, respectively.
Ablation Study. Finetuning Stages: In Table 3, we demonstrate the effectiveness of CALRec’s key components via extensive ablation study on $3$ data categories (the results on the other $3$ categories show the same trends so we skip them due to space constraints). First, it is obvious that our Stage I multi-category joint fine-tuning and Stage II category-specific fine-tuning are both effective, without which there will be an obvious drop in performance. Furthermore, it is worth mentioning that our model derived only with Stage I is able to recommend items in all $6$ categories: this is especially valuable and easy to deploy in real-world applications, due to its multi-tenancy nature. Although the multi-category trained model (variant B in Table 3) lags behind the full CALRec derived with both stages, it is already on par with variant A models (cf. Table 3)). We also compare CALRec and pretrained ‘off-the-shelf’ PaLM-2 (XXS) in Table 3; the results show that without any fine-tuning, PaLM-2 demonstrates poor performance on our sequential recommendation benchmarks.
Contrastive Objectives: In Table 4 we calculate the average scores on all $6$ categories for the setups with and without the auxiliary contrastive losses training objective. Basically, the auxiliary contrastive objective can result in circa $0.8\%-1.7\%$ gains. We also conduct dependent paired sample t-test (Ross and Willson, 2017) and report the p-values in the last column of Table 4: the gains are all statistically significant at $0.05$ level.
Model Size: We are also interested in how pretrained PaLM-2 without any fine-tuning on our sequential recommendation data performs in our tasks and we present results with three PaLM-2 models of three different model sizes (i.e., XXXS, XXS, and S) on ‘Scientific’ and ‘Instruments’ categories in Table 5. The input/output format is still the same: input data is from a single user. The results show that XXS and S models actually achieve similar performance and they both outperform the XXXS model.
6. Further Analyses and Discussion
W/O Dedup - Metric (opt) W/ Dedup - Metric (opt)
Dataset NDCG@10 Recall@1 Recall@10 MRR NDCG@10 Recall@1 Recall@10 MRR
Scientific 0.088 0.0355 0.1381 0.0745 0.0565 (-36%) 0.0142 (-60%) 0.1005 (-27%) 0.0461 (-38%)
Instruments 0.0545 0.0301 0.081 0.0485 0.0344 (-37%) 0.0040 (-87%) 0.0696 (-14%) 0.0263 (-46%)
Arts 0.0946 0.0511 0.1464 0.0811 0.0443 (-53%) 0.0067 (-87%) 0.0859 (-41%) 0.0344 (-58%)
Office 0.0757 0.0424 0.1095 0.0669 0.0391 (-48%) 0.0056 (-87%) 0.0735 (-33%) 0.0309 (-54%)
Games 0.05 0.0275 0.0755 0.0444 0.0211 (-58%) 0.0015 (-95%) 0.0442 (-41%) 0.0171 (-61%)
Pet 0.069 0.0318 0.1085 0.0578 0.0391 (-43%) 0.0094 (-70%) 0.0680 (-37%) 0.0317 (-45%)
Table 6. Data deduplication. Optimistic evaluations of LIR Baseline (always recommend last-item in the input sequence) on data from (Li et al., 2023b) (w/o dedup) and our data w/ dedup respectively.
Output Matches an Item (%) Hierarchical Match - Metric: (opt)/(pes) & Improv. vs BM25)
Dataset Format Error (%) All Four Attributes Title Only NDCG@10 Recall@1 Recall@10 MRR
Scientific 0.0% 82.3% 96.7% 0.0771 / 0.0725 (-2.1%) 0.0525 / 0.0404 (+3.2%) 0.1054 / 0.1054 (-5.9%) 0.0693 / 0.0632 (-5.2%)
Instruments 0.001% 97.6% 99.8% 0.0905 / 0.0901 (-0.4%) 0.0717 / 0.0705 (-0.1%) 0.1151 / 0.1151 (-0.6%) 0.0836 / 0.0830 (-3.2%)
Arts 0.01% 92.5% 98.9% 0.0854 / 0.0844 (-1.1%) 0.0636 / 0.0611 (+0.1%) 0.1113 / 0.1113 (-2.3%) 0.0777 / 0.0764 (-4.6%)
Office 0.0% 84.5% 94.7% 0.0926 / 0.0916 (-5.2%) 0.0736 / 0.0710 (-3.3%) 0.1127 / 0.1127 (-7.1%) 0.0864 / 0.0851 (-6.6%)
Games 0.05% 98.0% 99.0% 0.0582 / 0.0573 (-2.1%) 0.0276 / 0.0257 (-6.8%) 0.0973 / 0.0972 (-1.3%) 0.0467 / 0.0456 (-8.4%)
Pet 0.03% 97.3% 99.0% 0.0733 / 0.0690 (-0.4%) 0.0573 / 0.0468 (+0.6%) 0.0924 / 0.0922 (-1.3%) 0.0677 / 0.0620 (-2.8%)
Table 7. Hierarchical match results. The relative gains/drops on $\frac{\text{opt}+\text{pes}}{2}$ are also reported against BM25 retrieval.
Further Discussion on User Sequence Dedup. Now, we further delve deeper into user sequence dedup issues. Although in real world users can make repeated activities, we still posit that the high ratio of consecutive identical interactions in the raw Amazon Review data are not likely to reflect normal/real users’ behaviors (further details in Appendix). While we cannot conclude that data deduplication is ‘correct’ or ‘wrong’, we reckon that the ‘LIR’ baseline results are meaningful and should be reported for model comparison purposes. We further run the LIR Baseline on 1) data from Li et al. (2023b) without dedup and compare with its results on 2) our data with dedup. Although the data statistics are slightly different, the comparisons are still meaningful, since data sizes are both considerable and thus are representative of the full distribution. Table 6 shows the optimistic metrics of LIR Baseline that always recommends last item in the input sequence on data w/o dedup (left four columns) and data w/ dedup (right four columns). All metrics suffer sharp reductions due to the dedup process. Specifically, Recall@1 drops $60\%$-$95\%$ after removing consecutive identical events. We emphasize the necessity of sequence interaction deduplication as it is critical to reliably measuring model quality.
Does CALRec Rely on BM25? One challenge of using generative LLMs, especially for general-purpose LLMs not tuned with task-specific data, for recommendation is that there is no guarantee that the predicted item exists in the item corpus. Therefore we use BM25, a fuzzy lexical matching algorithm, to rank the items in the item corpus. However, the fine-tuned CALRec models can quickly master the item template format and remember items seen during training, making direct text-entity based retrieval possible without relying on fuzzy text matching methods. This motivates a simple and faster alternative to BM25.
First, we found that almost all ($>99.9\%$) finetuned model output honors the desired item text format as “Title: **. Category: **. Brand: **. Price: **.”. We then extract the four attribute entities: title, brand, category and price, and find that $82\%-98\%$ of model outputs can find an item in the item corpus that matches in all four attributes, and $95\%-99.8\%$ of title entities extracted from the model outputs exists in the item corpus. These statistics are summarized in Table 7. Based on these findings, we propose a heuristic hierarchical matching method. We first search exact matches in the item corpus for all four attributes extracted from model outputs; If no match is found, we search for the exact match of the first three attributes, then the first two, and finally only the title description. Similar to our main experiments, both optimistic and pessimistic metrics are calculated and results are also shown in Table 7: we find that our hierarchical matching method has marginally worse performance compared to the BM25-based retrieval but still outperforms other baselines models shown in Table 2.
Limitations and Future Work. Due to the fact that fine-tuned CALRec usually outputs exact item descriptions seen during training, a direct limitation ensues: CALRec barely cannot address cold-start scenarios where the ground-truth item is not seen in the training nor in the validation set. We leave addressing this limitation to future work.
Second, although CALRec outperforms other baselines in Recall@1, NDCG@10, and MRR@$\infty$ by large margin, its Recall@10 underperforms one of the text+ID combined baseline. We attribute this to the fact that temperature sampling, unlike embedding-based nearest neighbor search, is a quasi-greedy ranking method and may not yield the best top 10 predictions among the entire item corpus. Therefore, although CALRec performs well in metrics that emphasize the order of top predictions (like Recall@1 and MRR), its top-10 retrieval prediction needs more diversity.
Third, in real-world applications, when the size of item corpus is extremely large (e.g., billions of items), BM25 retrieval becomes extremely slow and even infeasible. In addition to the hierachical entity matching method discussed in the previous paragraph, we may extend CALRec to a retrieve-and-re-ranking model where any baseline sequential recommendation approach can be used to first derive, for example, top 10K recommendations focused on high recall. We can then make use of CALRec’s output texts to rerank these items which can not only leverage the strong capabilities of CALRec but also avoid doing BM25 retrieval on an extremely large corpus.
7. Conclusion
We proposed CALRec, a contrastive-learning-assisted two-stage training framework for sequential recommendation based on LLMs, where we run experiments with PaLM-2 LLM backbone. Our approach also features careful template design inspired from few-shot learning and a novel quasi-round-robin BM25 retrieval approach. Through comprehensive experiments on the Amazon Review dataset, we demonstrated the effectiveness of the proposed method supported by PaLM-2, where we outperform existing SotA models for sequential recommendation by a considerable margin in most evaluation metrics. We also presented a series of further analyses and revisited the user sequence duplication issue and existing evaluation metrics; we accordingly propose to adopt a training-free text-statistic based ‘Last Item Repeater’ reference baseline for model comparison purposes, and we strongly suggest reporting both optimistic and pessimistic metric scores for all text-based approaches.
Appendix A Examples of Data Duplication
As discussed in §4, the raw amazon review datasets contain repeated, non-distinguishable user-item interactions that share identical entity values across all attributes, including item id, user id, timestamp, item metadata, user rating, user review text, etc. An example in the industrial and scientific category is as follows:
"uid": "A3SFSFJZFI0OQN",
"asin": ["B00002NC3K", "B0001MSC84", "B0001MSC84", "B0001MSC84", "B0001MSC84"],
"overall": [3, 3, 3, 3, 3],
"reviewText": [
"Warped when exposed to the sun.",
"Warped when exposed to the sun.",
"Warped when exposed to the sun.",
"Warped when exposed to the sun.",
"Warped when exposed to the sun.",
"unixReviewTime": [1508544000, 1508544000, 1508544000, 1508544000, 1508544000],
"title": [
"Rubbermaid Commercial Products FG263100GRAY Rubbermaid Commercial Round Brute Container Lid, Gray, 32G",
"Rubbermaid Commercial BRUTE Heavy-Duty Round Waste/Utility Container with Venting Channels, 20-gallon, Gray (FG262000GRAY)",
"Rubbermaid Commercial BRUTE Heavy-Duty Round Waste/Utility Container with Venting Channels, 20-gallon, Gray (FG262000GRAY)",
"Rubbermaid Commercial BRUTE Heavy-Duty Round Waste/Utility Container with Venting Channels, 20-gallon, Gray (FG262000GRAY)",
"Rubbermaid Commercial BRUTE Heavy-Duty Round Waste/Utility Container with Venting Channels, 20-gallon, Gray (FG262000GRAY)",
In the example provided, all purchase-review events, except the initial one, are identical. This behavior generally deviates from normal user activities and may potentially confuse or mislead models into recommending repetitive items to achieve artificially inflated metrics. Although only a small percentage of the events are duplicates (Table 1), our analysis via the LIR baseline on raw and dedup datasets (Table 6) shows that those duplicates can have a significant impact on the behavior of the models because they provide models with an easy, oversimplified prediction path.
Appendix B Effectiveness of Contrastive alignment
Fig 2 shows the loss profiles during category-specific finetuning on the Scientific category without first multi-category joint fine-tuning. The left panel shows the model start overfitting in text generation task at $100$ epochs in the industrial and scientific category. The right panel shows that train, evaluation and test contrastive loss between prediction and target drops quickly from $6.2$ to $4.4$ and remains relatively stable. At batch size $512$, the initial $\mathcal{L}_{TT}\approx 6.2$ means the average ratio between $\exp(\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{i}^{T})/\tau_{c})$ and $\exp(\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{j}^{T})/\tau_{c}),j\neq i$ is about $1.04512\times\exp(-6.2)=$ 1.04 at the beginning of the training. Thus $\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{i}^{T})-\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{j}^{T})=\ln(1.04)\tau_{c}\approx 0.02\ll 1$, suggesting that without finetuning, the raw pretrained LLMs, although able to make some prediction (Tab 5), does not yield prediction embeddings that are more aligned with the ground truth items than random items. After finetuned and $\mathcal{L}_{TT}\approx 4.4$, the same ratio increases to $6.3512\times\exp(-4.4)=$ 6.3, corresponding to $\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{i}^{T})-\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{j}^{T})\approx 0.92\sim 1$, indicating that the predicted item embedding is $\sim 90^{\circ}$ more aligned to target embedding than a random item embedding. On the contrary, without contrastive alignment, the prediction-target contrastive loss $\mathcal{L}_{TT}$ only drops from $6.4$ to $6.0$ as the model learns to generate the correct next-item prediction as a pure text generation task. Fig 2 demonstrates that the contrastive loss can efficiently align the model prediction embedding and the target embedding.
Figure 2. Text generation cross-entropy loss ($\mathcal{L}_{NIG}$, left panel) and prediction-target contrastive loss ($\mathcal{L}_{TT}$, right panel) during Scientific-category-specific finetuning without multi-category joint fine-tuning. In both panels, the training curves (red-dotted) without contrastive alignment are also plotted for comparison.